Please write neatly. thank you

4. Solve the system X= 2
Answer #1

*** =(. :] Eigen values of A-1A-111=0 120d all 1-1 zodlo = (2-)2-1=0 → 4+d? 44-1=0 → 4-48 +3=0 dial 12=3 Eigen vector for d =ry DE IN-C: R₂ → R₂-Ri Ri Ricos) له و Let Vaat → Y = -t (:1-C2 : Eigenvalue 1: signvestor [:] Eigenvalne 3: eiguvector (:] ..